how old am i if 500 reduced by 3 times my get is 374

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)

Please check the problem before posting
how old am i if 500 reduced by 3 times my age is 374
I assume this is what you mean
let my age be x
500 reduced by 3 times my age is 374
500-3x = 374
subtract 500 from both sides
500-500 -3x = 500-374
-3x =-126
divide by -3
-3x/-3 = -126/-3
x= 42
your age is 42










Answer by ikleyn(52781)

.

Dear tutor @mananth

You produce perfect solutions, and I am very glad seeing it (!)

Can you please present them accurately by leaving no more than 4 empty lines after your post.


When you leave 10 - 20 empty lines after your post, it does not look good.

With 4 empty lines after the post, they will look much better.


Thank you (!)